OpenStudy (anonymous):

Pleasee help !! Multiply or divide 5x+10/ x+2 * 2x/4-10

OpenStudy (anonymous):

looks like multiply

OpenStudy (anonymous):

Hm. I think so to but I'm not sure how to solve it

OpenStudy (anonymous):

thats kinda hard to explain

OpenStudy (anonymous):

\[\frac{ 5x + 10}{x+2}\times \frac{ 2x }{ 4-10 }=\frac 5{ x + 2}{(x+2)}\times \frac{ 2x }{ -6}\] \[5\times \frac{ x }{ -3}=\frac{ 5x }{ -3}\]

OpenStudy (anonymous):

@dpasingh thank you so much , could you help me with another one?

OpenStudy (anonymous):

Yes, pls post the problem.

OpenStudy (anonymous):

3n^2-n/n^2 -1 divided by n^2/n+1

OpenStudy (anonymous):

Ok pls wait..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ 3n ^{2} -n}{ n ^{2} -1}\div \frac{ n ^{2} }{ n+1 }=n \frac{ 3n -1}{ (n -1)(n+1)}\times \frac{ n+1 }{n ^{2} }\] \[=\frac{ 3n -1}{ (n -1)}\times \frac{ 1 }{n }=\frac{ 3n -1}{ n(n -1)}\]

OpenStudy (anonymous):

thank youu !